the sum of the present ages of Vatha and Chris is 36. In 4 years time, the sum of their ages will equal twice Vatha's present age. How old are they now?
Answer by stanbon(75887)

the sum of the present ages of Vatha and Chris is 36. In 4 years time, the sum of their ages will equal twice Vatha's present age. How old are they now?
----
Equations:
v + c = 36
(v+4)+(c+4) = 2v
-----
Substitute and solve.
Cheers,
Stan H.